OpenStudy (itsbribro):

Mia deposits either $5 or $3.50 into her savings account. She made 14 deposits equaling a total of $59.50. How many times did she deposit $5 and how many times did she deposit $3.50?

OpenStudy (itsbribro):

Mia made ____ deposits of $5 and ____ deposits of $3.50.

OpenStudy (itsbribro):

@freckles could you help with this one please

OpenStudy (itsbribro):

@bubbles-are-cool.

OpenStudy (itsbribro):

@helpout3

OpenStudy (xapproachesinfinity):

let try the let call x number of deposits for $5 and y ......................................... for $3.5 total deposits x+y=14 5x number of deposit for $5 3.5y ....................... for $3.50 then the second equation is 5x+3.5y=59.50

OpenStudy (xapproachesinfinity):

substitute x with 14-y in the second equa to find x and then find y later

OpenStudy (xapproachesinfinity):

you know how to solve a system of two equation right? x+y=14 5x+3.5y=59.50

OpenStudy (itsbribro):

no i dont

OpenStudy (itsbribro):

:/